OpenStudy (anonymous):

Given: rectangle ABCD is similar to rectangle ZBXY. Enter answer as a simplified, improper fraction. If BZ = 3, BA = 8, BC = 12, then BX =

OpenStudy (anonymous):

this one I think I can do right

OpenStudy (anonymous):

\[\frac{8}{12}=\frac{3}{BX}\]\[ BX=\frac{12\times3}{8}\] BX=4.5

OpenStudy (anonymous):

u can check some one answer if I'm right